### v1

# TRec: Egocentric Action Recognition using 2D Point Tracks

Abstract
We present a novel approach for egocentric action recognition that leverages 2D point tracks as an additional motion cue. While most existing methods rely on RGB appearance, human pose estimation, or their combination, our work demonstrates that tracking randomly sampled image points across video frames can substantially improve recognition accuracy. Unlike prior approaches, we do not detect hands, objects, or interaction regions. Instead, we employ CoTracker to follow a set of randomly initialized points through each video and use the resulting trajectories, together with the corresponding image frames, as input to a Transformer-based recognition model. Surprisingly, our method achieves notable gains even when only the initial frame and its associated point tracks are provided, without incorporating the full video sequence. Experimental results confirm that integrating 2D point tracks consistently enhances performance compared to the same model trained without motion information, highlighting their potential as a lightweight yet effective representation for egocentric action understanding.
1 Introduction
The task of action recognition involves identifying and classifying human actions from video data. In the egocentric setting, videos are captured from a first-person perspective, typically using head-mounted cameras. Unlike exocentric (third-person) videos, which offer a stable, external view of actors performing generic activities such as sports or daily tasks, egocentric videos focus on fine-grained, domain-specific actions centered around hand–object interactions and the wearer’s visual attention. This perspective introduces unique challenges, including strong camera motion, frequent occlusions, and limited field of view, all of which complicate the modeling of both actor and object dynamics.
The prominent camera motion paired with the exceptional close and fast-moving background covering an extensive amount of the image pose an intricate challenge of modelling both actor and object dynamics. The interplay of intense camera motion combined with the rapidly moving and closely captured background dominating the visual field make it particularly challenging to model both actor and object dynamics. Many existing approaches rely on RGB-based features. While such methods have achieved strong results, they often depend on computationally expensive or error-prone semantic detection pipelines for hands, objects, or skeletons.
In this work, we investigate whether simple, semantically agnostic motion cues, specifically random 2D point tracks, can provide comparable or even superior information for recognizing actions in egocentric videos.
By tracking multiple points on hands, objects, and background regions, our method captures diverse motion signals, including head movement and object deformation. The difficulty of distinguishing object categories is reduced, since our method directly tracks the motion of the object in the scene. We generate these tracks using CoTracker. The classification pipeline uses the ResNet18 as backbone and a Transformer based module on top for classification. Surprisingly, our experiments show that a Transformer model trained with such point tracks, even when only the initial frame is used, can recognize actions more accurately than the RGB-only baseline using the whole video.
Our main contributions are as follows:
  • We propose a simple and efficient method for egocentric action recognition that integrates 2D point tracks with image features.
  • We demonstrate that these tracks capture meaningful motion patterns such as hand, object, and head movements without requiring explicit semantic detection.
  • We show consistent performance improvements over the RGB-only baseline on an egocentric benchmark.
2 Related Work
Trajectory-based methods model motion by tracking key points or regions across consecutive frames [18, 12, 20]. They have been explored in exocentric (third-person) action recognition, where stable camera viewpoints simplify trajectory estimation. To the best of our knowledge, no prior work has investigated the use of 2D point tracks for egocentric action recognition without filtering background points, where strong camera motion and frequent occlusions make trajectory-based modeling more challenging. Additionally, in contrast to these methods we utilize a relatively sparse set of points to track within the image. Wang et al. [18] introduced Improved Dense Trajectories, which compensate for camera motion using RANSAC-based estimation and employ a human detector to distinguish foreground motion from background motion. This approach established a strong pre-deep-learning baseline for video action recognition. Nguyen et al. [12] proposed Denser Trajectories of Anchor Points for Action Recognition, which anchor trajectories to more stable points to better capture fine-grained motion variations and increase robustness to camera movement.
Hand-object interaction modeling can improve egocentric action understanding. Several methods utilize some kind of hand-object detector and integrate this information into their model [3, 14, 13, 11]. Chatterjee et al. [3] expanded the vocabulary of egocentric actions by linking detected hands and objects with fine-grained action categories. Shiota et al. [14] explicitly modeled hand-object contact and object state transitions, while Peirone et al. [13] proposed a zone-aware recognition frameworks which learns environmental affordances to extract environment and action information seperately to train an action recognition model. Materzyńska et al. [11] proposed Something-Else, a compositional model that builds spatial-temporal interaction graphs between hands and objects, achieving strong performance on the Something-Something-v2 dataset. Although effective, these methods depend heavily on accurate hand and object detection, which can fail under occlusions or strong motion.
RGB-based approaches [15, 1, 9, 5, 16, 19] rely solely on raw visual appearance for action recognition, without incorporating explicit motion or geometric cues. Many of these high-performing models are trained on large-scale datasets, often leveraging extensive pretraining and carefully designed architectures. In contrast, we intentionally adopt a simple model design to clearly demonstrate the impact of incorporating point tracks. Girdhar et al. [5] propose the Anticipative Video Transformer, which models temporal dependencies and anticipates future actions in egocentric video sequences. Their results show that the model implicitly attends to image regions containing hands and objects, even without explicit supervision for their detection. Tong et al. [16] employ masked autoencoder pretraining to improve video representation learning, achieving strong performance on egocentric and other video benchmarks. Wang et al. [19] further extend pretrained video backbone models through an interactive prototype learning framework, enhancing active-object representations by leveraging actor motion cues within RGB inputs.
RGB-based approaches [15, 1, 9, 5, 16, 19] rely solely on raw visual appearance for action recognition, without incorporating explicit motion or geometric cues. Many of these high-performing models are trained on large-scale datasets, often leveraging extensive pretraining and carefully designed architectures. In contrast, we deliberately adopt a simple model design to isolate and clearly demonstrate the effectiveness of incorporating 2D point tracks. Unlike RGB-based features that primarily encode texture and color statistics, point trajectories provide explicit geometric and temporal information, capturing object and camera motion directly from the image sequence.
Girdhar et al. [5] propose the Anticipative Video Transformer, which models temporal dependencies in egocentric videos to predict upcoming actions. Their model implicitly attends to image regions containing hands and objects, and its performance improves with increased temporal context. However, the paper does not clarify whether motion is inferred implicitly through appearance changes between frames. In contrast, our approach introduces motion information explicitly by integrating 2D point tracks, enabling the model to reason directly about the displacement and dynamics of scene elements rather than relying on learned appearance differences. This explicit representation of motion complements the visual features and enhances interpretability in modeling fine-grained egocentric actions.
Tong et al. [16] employ masked autoencoder pretraining to improve video representation learning, achieving strong performance on egocentric and other video benchmarks. Wang et al. [19] further extend pretrained video backbone models through an interactive prototype learning framework, enhancing active-object representations by leveraging actor motion cues within RGB inputs.
Point tracking aims to estimate the trajectories of 2D points across video frames, providing fine-grained motion information that supports higher-level tasks such as action recognition and object interaction analysis. Recent advances in this area have focused on improving both tracking accuracy and computational efficiency. CoTracker [8] introduced a transformer-based architecture capable of jointly tracking multiple points across frames. Subsequent versions have further improved robustness, accuracy, and inference speed, establishing CoTracker as a state-of-the-art solution for general-purpose point tracking. In our work, we adopt CoTracker due to the public availability of pretrained weights, which allows for efficient and reliable integration of high-quality point trajectories into our framework. EgoPoints [4] proposed a dedicated benchmark for dense point tracking in egocentric videos. This work specifically addresses the challenges of egocentric settings, such as strong camera motion and frequent occlusions, and provides a comprehensive evaluation protocol for tracking performance. In addition, the authors improve upon CoTracker by adapting it to egocentric motion patterns and by providing a large-scale dataset for benchmarking.
3 Action Recognition using Point Tracks
We present TRec, a transformer-based model for egocentric action recognition that jointly leverages 2D point tracks and image representations. TRec tracks both background and foreground points throughout a video to capture rich motion cues without relying on explicit object or hand detections. The model employs a lightweight image encoder to extract spatial features from individual frames and a transformer architecture [17] to integrate temporal dependencies across frames. A set of randomly sampled 2D points is initialized within each video, covering both foreground and background regions, and their trajectories are tracked over time. The transformer receives as input the encoded image features together with these 2D point tracks, enabling it to reason jointly about appearance and motion information. A class token and multi-headed attention pooling are used to aggregate the transformer outputs, and the resulting representation is passed through an MLP-based action recognition head to predict the action depicted in the video.
We evaluate TRec on the Something-Something-v2 [6] dataset.
In the following, we first describe the dataset curation process, including sampling strategies and preprocessing steps used to generate point trajectories. Next, we present the point tracking procedure and explain how trajectories are integrated with image features. We then detail the overall model architecture, including the image encoder, transformer design, and the action recognition head. Finally, we outline the training setup and evaluation protocol used to assess TRec on egocentric benchmarks.
3.1 Dataset Curation and Preprocessing
For training and evaluation, we use the Something-Something-v2 dataset [6].We employ CoTracker [8] to obtain 2D point trajectories across video frames.
Something-Something-v2. The Something-Something-v2 dataset [6] contains short video clips depicting human-object interactions recorded in a wide variety of everyday environments. It comprises 174 action templates describing generic activities of “someone doing something,” such as pushing, pulling, or placing objects. It offers greater diversity in both visual context and action types, as well as a substantially large number of annotated samples, making it a valuable benchmark for training data-driven models. Because the camera remains largely stationary across most recordings, the dataset provides a stable visual setting that facilitates the analysis of object and hand dynamics without the confounding effects of egocentric motion. We use the dataset in its original form without any additional preprocessing or modifications.
Point Track Generation. We curate 2D point trajectories using CoTracker3 [7]. For each action sequence, we randomly sample 900 points in a covering the entire image, including both foreground and background regions. Video frames are sampled at 30 fps to obtain smooth trajectories. The resulting point coordinates are normalized following the procedure described in [2]. For each video we track a set of 900 points $p^{i}=(x^{i},y^{i})\in\mathbf{R}^{2}$. Here, $i\in[0,P]$ denotes the $i^{\text{th}}$ point with $P$ being the total number of points. The point tracks are denoted as $p_{t}^{i}\in\mathbf{R}^{T\times N\times 2}$ with $t\in[0,T]$ refering to the frame index, with $T$ denoting the number of frames in the video.
(a) Hitting steel glass with screw driver
(b) Lifting a lighter and dropping it
(c) moving scissors and screwdriver closer
Figure 1: These images show the tracked trajectory of 2D points in videos of the Something-Something dataset.
3.2 Network Architecture
An overview of the proposed network architecture is shown in Figure 2. The input to the network is a video $v\in\mathbf{R}^{T\times H\times W\times 3}$ where $H,W$ correspond to the image height and width, respectively. Additionally, the model receives 2D point tracks $p^{i}_{t}$ extracted from the video. The architecture consists of an image encoder $I(v_{t})$ with $v_{t}\in\mathbf{R}^{h\times w\times 3}$ as backbone and a transformer module $G:(i,p)\rightarrow g$ which takes the image features $i$ extracted by the encoder and the point tracks $t$ as input. The output $g$ of the transformer is pooled by an additional multi-head attention layer $M:g\rightarrow o$ which aggregates information from all tokens. Finally, the class token is passed through an MLP-based classification head to predict the action label.
Figure 2: Network Architecture
3.3 Action Recognition
The transformer model receives as input the image features extracted by the encoder together with the corresponding 2D point tracks. This joint representation allows the model to integrate appearance and motion information in a unified latent space. Each frame contributes both spatial features from the image encoder and geometric motion cues from the tracked points. By processing these modalities jointly, the transformer learns to associate appearance changes with underlying motion dynamics, enabling robust recognition of fine-grained actions in egocentric settings.
Within the transformer, spatio-temporal dependencies are modeled through self-attention across both frame and point dimensions. This mechanism allows the network to focus on motion patterns that are most informative for distinguishing between actions such as subtle object deformations or coordinated hand movements without requiring explicit hand or object detection. The multi-head attention pooling layer then aggregates the sequence of token embeddings into a single representation guided by the class token. The resulting feature vector encodes both global context and localized motion cues relevant to the observed activity.
Finally, the pooled class token is passed through a lightweight multi-layer perceptron (MLP) classification head to produce the final action prediction. During training, the model is optimized using a standard cross-entropy loss over the action classes defined by each dataset. This simple yet effective design allows TRec to capture meaningful motion representations directly from tracked points and image features, achieving competitive performance while maintaining low computational complexity.
4 Experiments
We conduct a series of experiments to evaluate the effectiveness of the proposed model under various configurations. All models are implemented using a ResNet18 backbone combined with a Transformer architecture. The Transformer module consists of six attention layers with four heads each and an intermediate dimension of 2048. Before being fed into the Transformer, the 2D point tracks $p$ are normalized and reshaped to $(B,T\times 2,P)$, after which they are passed through an MLP to match the input dimensionality of the Transformer. The number of frames per video is set to 8 if not stated otherwise. The number of point tracks is randomized between 200 and 400 points for each mini-batch.
We compare two model variants: (1) a baseline using only RGB input, (2) a trackpoint model that incorporates 2D point trajectories, and
The baseline receives only the video frames as input. Unless otherwise stated, training is performed with a batch size of 32, AdamW optimizer [10], a learning rate of $1\times 10^{-4}$ for the Transformer, and $1\times 10^{-5}$ for the backbone network. We use a cosine warm-restart learning rate scheduler with 5000 warm-up iterations and train for 20 epochs on an NVIDIA A40 GPU. Input frames are resized to $256\times 256$ and augmented using random cropping, horizontal flipping, Gaussian blur, and color jitter. The tracks are agumented according to the augmentation of their corresponding image. The observation horizon is set to $8$ frames. Out of the 900 tracked points generated per video, 400 points are randomly selected and provided to the network to promote diversity. The random number generators are seeded in all tests, thus in each experiment the model receives the same point tracks.
4.1 Track vs no Track
The first experiment compares the performance of TRec with a baseline model that shares the same architecture and training parameters. The results, summarized in Table 1, demonstrate that TRec achieves substantially higher accuracy than the baseline. This improvement indicates that the model effectively leverages the 2D point tracks to capture motion dynamics and accurately infer the actions occurring in the video.
Table 1: Track vs No-Track
Model Top-1 Top-5
TRec $61.10\%\pm 8.66\%$ $83.95\%\pm 6.62\%$
baseline $30.27\%\pm 8.05\%$ $53.24\%\pm 8.75\%$
These models are trained with 8 frames with their corresponding point tracks as input.
4.2 Number of points
To assess the influence of the number of tracked points on action recognition performance, we conduct an ablation study by varying the number of input points provided to the model. This experiment investigates how sensitive TRec is to the density of motion information encoded in the point trajectories. The quantitative results are summarized in Table 2, while Figure 3 illustrates how performance changes with different numbers of points. The model maintains stable accuracy when using 50 or more tracked points, with diminishing performance gains beyond 100 points. However, when fewer than 25 point tracks are used, the model’s accuracy drops noticeably, indicating that a minimal number of trajectories is required to capture sufficient motion cues for reliable action recognition.
Table 2: Ablation on the number of tracked points. Performance is reported on the Something-Something-v2 datasets.
Number of Points Top-1 Top-5
[…]
100 $60.61\%\pm 8.58\%$ $82.66\%\pm 6.67\%$
50 $59.73\%\pm 8.71\%$ $82.11\%\pm 6.72\%$
25 $57.75\%\pm 8.72\%$ $80.60\%\pm 6.93\%$
15 $54.98\%\pm 8.68\%$ $78.24\%\pm 7.25\%$
5 $45.19\%\pm 8.85\%$ $69.10\%\pm 8.20\%$
1 $24.87\%\pm 7.68\%$ $46.53\%\pm 8.82\%$
Figure 3: Ablation on the number of point tracks used as input to the model.
4.3 Background Motion Filtering Using Kernel Density Estimation
To investigate the influence of background motion on action recognition, we analyze the effect of filtering point tracks using a Kernel Density Estimation (KDE) approach. Our main hypothesis is that training the model on filtered points, which primarily represent hands and manipulated objects, will lead to slightly lower performance compared to using all points, as background motion provides valuable contextual cues for recognizing actions.
Filtering is performed by applying KDE to the motion vectors of the tracked points, as illustrated in Figure 5. Since most points in the image belong to the background and therefore share similar motion patterns, the KDE identifies these high-density motion clusters corresponding to background movement. By excluding these clusters, we retain mainly the low-density points associated with hands and active objects. We then train a model using only the filtered points and compare its performance to a model trained on the full set of tracks that includes both foreground and background motion. The model trained only on filtered points is denoted here as FilterTRec. The quantitative results of this comparison are presented in Table 3.
The standard TRec model achieves a Top-1 accuracy of $61.10\%$. However, when the input points are filtered using KDE, performance drops notably to $52.46\%$. This decline indicates that the model effectively utilizes background motion cues and partially relies on them for action prediction. To further validate this observation, we trained a variant of our model exclusively on the filtered points. As shown in Table 3, the performance decreases even further (Top-1 accuracy of $36.34\%$), confirming that background motion contributes significantly to the model’s recognition capability.
Table 3: Model evaluation with Kernel Density Function
Model SSv2
Top-1 Top-5
vanilla TRec $61.10\%$ $83.95\%$
vanilla TRec + KDE $52.46\%$% $74.04\%$
FilterTRec + KDE $36.34\%$ $60.57\%$
Figure 4: Points without filtering
Figure 5: Points filtered using KDE
4.4 Single-Image Evaluation
To further analyze the contribution of the tracked points, we train and evaluate the model using only a single image while retaining the complete set of point tracks from the corresponding video. This experiment isolates the influence of motion trajectories from visual appearance and emphasizes the model’s ability to infer actions purely based on the temporal evolution of the tracked points.
In this setup, only the initial image of each video is provided as input during training and testing. Because this frame may not contain visible hands or objects, the model must rely entirely on the tracked motion to infer hand movements and object interactions. Moreover, the Something-Something-v2 dataset includes several action categories that can only be recognized by observing the final frames of a video, such as pretending to do something or revealing something. Since the model cannot predict these actions without the last frame, we evaluate the approach both on the unmodified dataset and on a version with these ambiguous categories removed, specifically those containing the terms pretend, uncover, reveal, fail in their annotations.
As shown in Table 4, the model achieves higher performance than the RGB-only baseline, demonstrating that the learned motion representations from the point tracks provide strong cues for action recognition even when only a single frame is available. As input we only provide the initial image. That image might not contain the hands, thus the model has to extract the hand motions from the point tracks.
Table 4: Single Image Evaluation
Model Top-1 Top-5
TRec 44.7% 69.12%
baseline 37.5% 62.41%
Performance on the Something-Something-v2 dataset using initial image as input and additionaly the point tracks for TRec.
5 Discussion
The experimental results demonstrate that incorporating 2D point tracks provides a clear advantage for egocentric action recognition compared to purely RGB-based methods. By explicitly encoding motion cues through tracked points, our model captures spatial-temporal dynamics that RGB alone cannot represent. This finding supports our initial hypothesis that motion trajectories offer a lightweight and interpretable alternative to computationally demanding approaches such as optical flow or hand-object detection.
The ablation on the number of tracked points indicates that the model is robust to variations in track density. Performance remains stable with as few as 50 points, and improvements beyond 100 points become marginal. This suggests that the model efficiently extracts the most relevant motion information without requiring dense tracking, offering a favorable trade-off between computational cost and accuracy. However, performance degrades substantially when fewer than 25 points are used, implying that a minimal level of motion diversity is necessary for reliable recognition.
The single-image experiment further highlights the discriminative power of the point trajectories. Even when only the first frame of a video is available, the model achieves higher accuracy than the RGB-only baseline, demonstrating that temporal motion encoded by the tracks carries sufficient information to infer the underlying action. This result also underlines the interpretability of our representation where the model can reason about hand and object movement even if they are not visible in the initial frame.
Finally, the kernel density filtering experiment reveals that background motion contributes significantly to action recognition in egocentric videos. Removing background points leads to a noticeable drop in accuracy, indicating that the model uses global motion patterns such as camera or head movements as contextual signals for understanding actions. This aligns with the egocentric nature of the data, where background dynamics often correlate with the actor’s motion and intention.
Overall, these findings confirm that point-track representations provide an effective and computationally efficient way to model motion in egocentric videos. Future work could focus on refining the integration of point-based and semantic cues, dynamically weighting background and foreground contributions, and developing improved approaches for collecting and maintaining stable track points in egocentric scenarios with substantial head movement, where existing tracking methods often degrade in performance.